# Data-Driven Probabilistic Air-Sea Flux Parameterization

### v2

@@ Acknowledgements.
This research received support through the National Science Foundation Science and Technology Center, Learning the Earth with Artificial Intelligence and Physics, LEAP (Grant number 2019625). We thank Julia Simpson, Pierre Gentine, and Bia Villas Bôas for helpful discussions and the anonymous reviewers for their useful feedback. Aakash Sane, Qing Li, and Stéphane Popinet provided valuable advice regarding the GOTM implementation. We acknowledge the continued efforts from NOAA PSL lab (including Ludovic Bariteau, Chris Fairall, and many others) in obtaining and publishing the EC datasets. The statements, findings, conclusions, and recommendations are those of the author(s) and do not necessarily reflect the views of the National Oceanic and Atmospheric Administration, or the U.S. Department of Commerce. This research was also supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.

### v3

@@ Acknowledgements.
This research received support through the NSF STC, Learning the Earth with Artificial Intelligence and Physics, LEAP (Grant number 2019625). We thank Julia Simpson, Pierre Gentine, and Bia Villas Bôas for helpful discussions and the anonymous reviewers for their useful feedback. Aakash Sane, Qing Li, and Stéphane Popinet provided valuable advice regarding the GOTM implementation. We acknowledge the continued efforts from NOAA PSL lab (including Ludovic Bariteau, Chris Fairall, and many others) in obtaining and publishing the EC datasets. The statements, findings, conclusions, and recommendations are those of the author(s) and do not necessarily reflect the views of the National Oceanic and Atmospheric Administration, or the U.S. Department of Commerce. This research was also supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.